OpenStudy (anonymous):

Use the table below to evaluate d/dx {f{g(2x)}} at x = 1. x 1 2 3 4 f(x) 6 1 8 2 f ′(x) 6 1 8 2 g(x) 1 4 4 3 g ′(x) 9 5 5 –4

OpenStudy (freckles):

do you know chain rule?

OpenStudy (freckles):

\[\frac{d}{dx}(f(g(u)))=\frac{du}{dx} \cdot g'(u) \cdot f'(g(u))\]

OpenStudy (anonymous):

ahhhh ok

OpenStudy (freckles):

I guess you don't need any help?

OpenStudy (freckles):

you can show me what you have after applying chain rule

OpenStudy (anonymous):

i mean i dont know how to do it

OpenStudy (freckles):

well i told you what chain rule is... the only difference between your expression and my expression is that my u is your 2x

OpenStudy (freckles):

replace u with 2x

OpenStudy (freckles):

and let me know what you have after that

OpenStudy (anonymous):

1

OpenStudy (freckles):

\[\frac{d}{dx}(f(g(u)))=\frac{du}{dx} \cdot g'(u) \cdot f'(g(u)) \] did you replace u with 2x?

OpenStudy (freckles):

\[\frac{d(2x)}{dx} g'(2x) f'(g(2x))\] find the derivative of 2x there at the beginning factor then replace all the x's with 1's

OpenStudy (freckles):

then use your chart to evaluate